OpenStudy (anonymous):

sqrt(34-24*sqrt(2)) * (4+ 3*sqrt(2))

OpenStudy (anonymous):

noone

OpenStudy (anonymous):

aneone

OpenStudy (anonymous):

\[\sqrt{34 -24\sqrt{2}} * (4+3\times \sqrt{2})\]

OpenStudy (bahrom7893):

hmmm let me try this on paper..

OpenStudy (anonymous):

ya

OpenStudy (bahrom7893):

do u just want the approximation?